### v2

# A Comparative Study of Image Restoration Networks for General Backbone Network Design

Abstract
Despite the significant progress made by deep models in various image restoration tasks, existing image restoration networks still face challenges in terms of task generality. An intuitive manifestation is that networks which excel in certain tasks often fail to deliver satisfactory results in others. To illustrate this point, we select five representative networks and conduct a comparative study on five classic image restoration tasks. First, we provide a detailed explanation of the characteristics of different image restoration tasks and backbone networks. Following this, we present the benchmark results and analyze the reasons behind the performance disparity of different models across various tasks. Drawing from this comparative study, we propose that a general image restoration backbone network needs to meet the functional requirements of diverse tasks. Based on this principle, we design a new general image restoration backbone network, X-Restormer. Extensive experiments demonstrate that X-Restormer possesses good task generality and achieves state-of-the-art performance across a variety of tasks.
Abstract
In this supplementary material, we present additional experiments and results to complement the main manuscript. First, we illustrate the importance of a general backbone network for multi-task optimization on all-in-one restoration. Then, we complement details about schematic diagrams of mentioned three kinds of architectures. Next, we provide complete benchmark results for the five image restoration tasks and present visual results of X-Restormer with other benchmark networks. After that, we conduct comprehensive ablation study to further verify the effectiveness of our X-Restormer. Finally, we provide a model complexity comparison of our X-Restormer with Restormer.
1 Introduction
Image restoration aims to generate high-quality images from degraded images. In recent years, deep learning has achieved great success in this field, with numerous networks being proposed to address various image restoration tasks. Initially, networks are primarily designed to solve specific restoration tasks and are typically validated only on selected tasks. As deep learning techniques have continued to evolve, there has been an increasing focus on the development of general-purpose networks that can be applied to a broad range of tasks. This trend is particularly evident in the high-level vision field, where new backbone networks are being designed to support multiple tasks [17, 30], including classification, detection and segmentation. For image restoration, although more and more restoration backbones can handle multiple restoration tasks, their task generality is still limited, as illustrated in Figure 2. For instance, SwinIR [25] achieves state-of-the-art performance on image super-resolution (SR) but falls short on image deblurring and dehazing. Conversely, Restormer [50] performs exceptionally well on image dehazing and deraining but is less effective on image SR. This discrepancy can be attributed to the fact that the characteristics of image degradation vary across different image restoration tasks. While all image restoration tasks involve mapping degraded images to clean images, the requirements for the capability of backbone networks differ depending on specific tasks.
Figure 1: Relative performance difference of different backbone networks on five image restoration tasks [^2] We set the minimum average performance of the networks on test sets in Table 2 for the task (i) as the lower bound $P^{(i)}_{lower}$, and the performance of X-Restormer as the upper bound $P^{(i)}_{upper}$. The ordinate of each point with performance $P^{(i)}$ is calculated as $(P^{(i)}-P^{(i)}_{lower})/P^{(i)}_{upper}$.. The existing representative networks exhibit diverse performance on these tasks, while our method presents superior task generality.
Designing a general image restoration backbone network presents a significant challenge. However, the development of such a network holds considerable value, as it has the potential to greatly reduce costs associated with research and application. To achieve this goal, we first conduct a comparative study of mainstream backbone networks on the representative tasks, including image SR, denoising, deblurring, deraining and dehazing. These five tasks are chosen due to the distinct characteristics of their degradation. Five representative backbone networks are selected in the study, including MPRNet [49], Uformer [44], SwinIR [25], Restormer [50] and NAFNet [6]. These five networks encompass classic architectures such as U-shape architecture, plain residual-in-residual architecture and multi-stage progressive architecture. They also employ several common operators, including convolution, spatial self-attention and transposed self-attention [50]. We benchmark the five representative methods on the selected five tasks. The experimental results cleaårly reflect the performance disparity of different backbone networks on different tasks. We then conduct a detailed analysis of the characteristics of these tasks and these backbone networks to explain the reasons behind the performance differences. Based on the comparative study, we believe that a general backbone network must be highly comprehensive in terms of functionality that meets the diverse needs of various tasks.
It is noteworthy that Restormer stands out in the comparative study, ranking within the top two across all five tasks. This superior performance can be attributed to several key designs. First, Restormer’s U-shape architecture allows it to process large-size inputs. Then, the network employs transposed self-attention, utilizing channel-wise feature maps as tokens, which enables mapping with a global receptive field. Additionally, the incorporation of numerous depth-wise convolutions activates the spatial information interaction. From a functional perspective, Restormer integrates the capabilities of the other compared networks, thereby exhibiting commendable task generality. However, the spatial mapping ability of Restormer appears to be somewhat deficient, as indicated by its quantitatively and qualitatively subpar performance in comparison to SwinIR for SR.
This inferiority is hypothesized to originate from the inherent challenge of detail reconstruction posed by the U-shape architecture, coupled with the relatively weak spatial mapping capability of depth-wise convolution, particularly when compared to spatial self-attention (i.e., window-based self-attention in SwinIR). To address this limitation, a plausible solution is the introduction of spatial self-attention into Restormer. To validate the effectiveness of our design, we alternately replace half of the transposed self-attention block with the overlapping cross-attention block [8] to build a new network, X-Restormer. This simple modification significantly enhances the performance of Restormer without increasing the number of parameters. Experiments reveal that X-Restormer obtains state-of-the-art performance on all five tasks, exhibiting the best task generality.
Our main contributions can be summarized as follows:
  • $\bullet$
    We conduct a comparative study by constructing an image restoration benchmark, highlighting the challenges faced by existing backbone networks in task generality.
  • $\bullet$
    Based on the benchmark results, we perform a detailed analysis of the characteristics of different degradations and different networks. We emphasize that the general backbone network design must meet the functional requirements of diverse restoration tasks.
  • $\bullet$
    By further enhancing the spatial mapping ability of Restormer, we design a preliminary general backbone network, X-Restormer. Without additional parameters, X-Restormer achieves significant performance improvement and exhibits superior task generality.
2 Related Work
Image restoration networks.
Since SRCNN [12] first introduces deep learning to image SR and obtains superior performance over conventional methods, numerous deep networks have been proposed for various image restoration tasks such as image SR [26, 60, 8], denoising [54, 55, 44], deblurring [1, 5], deraining [48, 4, 29] and dehazing [46, 41, 40]. In earlier years, most deep networks are designed for specific image restoration tasks. For instance, [59] uses the channel attention mechanism to construct a CNN-based SR network. [56] builds a denoising network by inputting multiple downsampled sub-images and a noise level map. [15] utilizes residual connections for a deraining network. [3] estimates the transmission rate and inverts the atmospheric scattering model using a CNN to achieve image dehazing. [34] designs a multi-scale network for image deblurring.
Figure 2: Selected five image restoration tasks. The degradation in each task varies significantly.
Recently, more and more methods have been developed to tackle multiple image restoration tasks. [49] builds a multi-stage CNN for deraining, deblurring and denoising. [44] designs a U-shape Transformer for deraining, deblurring and denoising. [25] implements a Swin Transformer-based network that achieves state-of-the-art performance on SR, denoising and compression artifact reduction. [50] proposes a novel transposed self-attention to build a U-shape network for deraining, deblurring and denoising. [6] constructs a U-shaped CNN for denoising and deblurring.
Difference from the previous network design research.
While previous works have proposed networks that excel in various image restoration tasks, their primary focus is on constructing stronger networks to achieve performance breakthroughs on specific tasks. In contrast, this work pays more attention to the task generality of the backbone network, possessing a vision different from previous works. More specifically, our objective is to explore the design principles and directions of general image restoration networks. We are not seeking to create powerful networks for peak performance on a single or some specific tasks, but rather to ensure satisfactory performance across a diverse range of tasks. Regarding the concrete implementation, we do not intend to construct complex network architectures or modules. Our preference, rather, is to enhance task generality through the use of the simplest methodology available.
There are concurrent works that adopt similar ideas for specific image restoration tasks. DAT [9] combines spatial-window self-attention and channel-wise self-attention to handle image SR. IPT-V2 [24] designs a spatial-channel Transformer block to build a denoising network and obtains the winner award in the NTIRE 2023 image denoising challenge [24]. However, the motivation and specific network implementation of our work are distinct from these studies.
3 Image Restoration Benchmark
In this section, we first briefly introduce several image restoration tasks, each with its own representative degradation characteristics. Subsequently, we classify mainstream image restoration networks based on two key aspects: architecture and core operator. On this basis, we select five representative networks and conduct a benchmark experiment across five different tasks. We describe the experimental setup and explain its rationality. Finally, we present the benchmark results and conduct a detailed analysis of them.
3.1 Overview of Image Restoration Tasks
We select five representative tasks for the benchmark experiments. These tasks, exemplified in Figure 2, are chosen based on two primary reasons. First, they are very common restoration tasks. Second, and more importantly, the degradation characteristics of these tasks vary significantly. As such, they can provide a robust means of evaluating the task generality of an image restoration backbone network.
Let $I_{GT}$ denote the ground truth image and $I_{LQ}$ denote the degraded image, where $I_{GT}\in\mathbb{R}^{H\times W\times 3}$. The degradation model of classic image SR can be represented as:
$I_{LQ}=(I_{GT}\otimes k)\downarrow_{s},$ (1)
where $I_{LQ}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times 3}$ represents the low-resolution image. $k$ denotes the bicubic downsampling kernel and $\downarrow_{s}$ represents the downscaling factor. This degradation is highly correlated to local information and leads to a significant loss of high-frequency information. Thus, SR networks emphasize strong spatial information interaction capability to reconstruct as many details as possible.
The degradation model of image denoising can be denoted as:
$I_{LQ}=I_{GT}+n,$ (2)
where $n\in\mathbb{R}^{H\times W\times 3}$ represents the noise map. For Gaussian denoising, noise values are content-independent. The downsampling-upsampling process of U-shape architecture inherently aids noise removal. Besides, strong spatial information interaction capability can also enhance high-frequency content reconstruction for denoising networks.
The degradation model of image deblurring (for motion deblurring) can be denoted as:
$I_{LQ}=\sum_{t}(f_{motion}^{t}(I_{GT})),$ (3)
where $f_{motion}^{t}(\cdot)$ represents the motion function under different continuous exposure times. This degradation is related to the global motion offset of the image. Therefore, the ability to utilize long-range information and even global information is important for deblurring networks.
The degradation model of image deraining can be simply denoted as:
$I_{LQ}=I_{GT}+R,$ (4)
where $R$ denotes the additive rain streak, simulated by the physics models, such as [23, 28]. The difference between this degradation and Gaussian noise is that the added $R$ is not evenly distributed on the image and has a correlation with the image content.
Figure 3: The core operators in image restoration networks.
The degradation model of image dehazing, based on the atmospheric scattering model, can be denoted as:
$I_{LQ}=I_{GT}*t(I_{GT})+A(1-t(I_{GT})),$ (5)
where $t(\cdot)$ represents the transmission function and $t(I_{GT})$ is associated with the distance from the scene point to the camera. This degradation is intrinsically linked to the depth information within the image. Consequently, the incorporation of global information is important for the effective functioning of dehazing networks.
3.2 Characteristics of Typical Backbone Networks
The architectures of mainstream image restoration networks can be broadly classified into three categories: U-shape encoder-decoder, plain residual-in-residual and multi-stage progressive. Schematic diagrams of these architectures are provided in Supp. The U-shaped encoder-decoder architecture performs down-sampling and up-sampling operations on features, enabling networks to handle features of varying scales. This architecture allows networks to accept large-size input, and the effective receptive field of the network expands rapidly with down-sampling. Typical U-shape networks include Uformer [44], Restormer [50]. The multi-stage architecture divides the entire network into several sub-networks and progressively processes features, which are primarily used for image deraining and deblurring. Common networks based on this architecture include MPRNet [49] and HINet [5]. The plain residual-in-residual architecture is composed of several residual groups, each of which consists of several residual blocks. This architecture maintains the original size when processing features, which is favorable for the reconstruction of high-frequency information, but it comes at a high computational cost. Typical networks include RCAN [59] and SwinIR [25].
The core operators for constructing an image restoration network can be mainly categorized into three types: convolution, spatial self-attention and transposed self-attention. These operators are shown in Figure 3. The convolution calculates a fixed-size filter and processes the entire feature map through a sliding window, which is the major component of many networks, such as RDN [62]. Spatial self-attention is typically implemented as window self-attention in image restoration tasks. It calculates the attention matrix within a fixed window size, generating content-aware weights that are functionally similar to a large kernel dynamic filter. This operator has strong local fitting ability and shows superior advantages on SR and denoising [7]. Transposed self-attention treats the entire feature of each channel as a token to calculate the attention matrix on the channel dimension. This operator directly deals with global features, and when combined with depth-wise convolution, it shows remarkable performance in multiple restoration tasks [50].
The selected five representative backbone networks for the benchmark experiment encompass the abovementioned architectures and core operators, as presented in Table 1.
Table 1: Architectures and core operators of the five selected backbone networks for the image restoration benchmark.
Network Architecture Core operator
MPRNet Multi-Stage Convolution
Uformer U-Shape Spatial self-attention
SwinIR Plain residual-in-residual Spatial self-attention
Restormer U-Shape Transposed self-attention
NAFNet U-Shape Convolution
3.3 Experimental Settings
For image SR, we conduct experiments on upscaling factor ×$\times 4$. We use the DF2K dataset (the same as SwinIR [25]) to train models. Low-resolution images are generated from the ground truth images using bicubic downsampling in MATLAB. For U-shape networks, we first up-sample the input low-resolution images through bilinear interpolation. The performance is reported on the Y channel. For denoising, we adopt the DFWB dataset for training. Noisy images are generated by adding Gaussian noise with a noise level of 50. For deblurring, we use the motion deblurring dataset GoPro [34] to train the models. For deraining, we conduct experiments using the synthetic rain dataset Rain13K and calculate the performance on the Y channel, following Restormer [50]. For dehazing, we use the indoor training set (ITS) of the RESIDE dataset [21], the same as [40].
To maximize the capability of these networks, we use the official codes and training configurations provided by different methods to train the models [^3] We have also tried to train all networks with a unified configuration, but find it unreasonable. The performance of networks may vary greatly with different training configurations and optimization strategies, making it difficult to determine a fair unified setting. . Note that all models are trained without using any pre-training strategy (e.g., ×$\times 2$ pre-training for SR) or special tricks (e.g., EMA in SwinIR and TLC in NAFNet) for fair comparison. In addition, we find that different methods may not use exactly the same test sets and the same metrics calculation in their papers to report performance. Therefore, we retest all models based on exactly the same data and calculate PSNR/SSIM using the popular open-source toolbox BasicSR [43].
Table 2: Quantitative results (PSNR/SSIM) of the benchmark. The best and second-best performance results are in bold and underline.
Method SR Denoising Deblurring Deraining Dehazing
Set14 Urban100 CBSD68 Urban100 GoPro HIDE Test100 Rain100H SOTS Indoor
MPRNet 28.90/0.7889 26.88/0.8081 28.48/0.8087 29.71/0.8847 32.66/0.9392 30.96/0.9188 30.29/0.8983 30.43/0.8905 40.34/0.9936
SwinIR 29.07/0.7944 27.47/0.8258 28.56/0.8118 29.88/0.8861 31.66/0.9209 29.41/0.8957 30.05/0.8996 30.45/0.8952 29.14/0.9676
Uformer 27.14/0.7398 25.60/0.7651 28.55/0.8130 29.98/0.8900 33.05/0.9417 30.89/0.9203 27.93/0.8910 24.06/0.8454 33.58/0.9856
Restormer 29.06/0.7934 27.32/0.8199 28.60/0.8130 30.02/0.8898 32.92/0.9398 31.22/0.9226 32.03/0.9237 31.48/0.9054 41.87/0.9938
NAFNet 29.03/0.7919 27.00/0.8112 28.52/0.8098 29.65/0.8840 33.08/0.9419 31.22/0.9237 30.33/0.9097 32.83/0.9135 38.97/0.9928
Figure 4: Visual comparison of Restormer and SwinIR. SwinIR reconstructs significantly more details than Restormer.
3.4 Benchmark Results
We present the quantitative results for each task on benchmark test sets in Table 2. (Due to space constraints, complete results are provided in Supp.) Several important observations can be made from the benchmark results: 1) Different networks exhibit varying performance on different tasks. For instance, SwinIR performs best on SR but worst on deblurring and dehazing. Uformer excels on denoising and deblurring but performs poorly on deraining and SR. 2) Networks with U-shape and multi-stage architectures demonstrate clear advantages on deblurring and dehazing. 3) MPRNet and NAFNet, which are mainly based on convolution operators, exhibit moderate performance across all tasks without outstanding results. 4) SwinIR, which employs plain architecture and spatial self-attention operators, outperforms other networks by a significant margin on SR. 5) The overall performance of Restormer is outstanding. Except for consistently being weaker than SwinIR on SR, it obtains considerable performance on almost all other tasks.
3.5 Analysis
We explain the above observations by analyzing the characteristics of different tasks and backbone networks.
The degradation of SR lies in the compression of local information, resulting in a large loss of high-frequency details. Therefore, SR networks often require strong spatial information interaction capability, or even generative capability. The U-shape architecture, which incorporates multiple downsampling operations, intuitively escalates the difficulty of detail reconstruction. In contrast, the plain architecture that maintains feature sizes benefits SR. Besides, window self-attention has demonstrated a superior local fitting ability than convolution [8]. As a result, SwinIR, which is based on a plain structure and employs spatial self-attention operators, exhibits a distinct advantage on SR.
Denoising entails smoothing the image to eliminate high-frequency noise and integrating low-frequency information to reconstruct a clear image. It is conjectured that the high performance of Restormer on denoising can be attributed to its ability to better smooth noise through channel-based processing, akin to operating in the frequency domain. Conversely, SwinIR performs well due to the robust reconstruction ability of its local self-attention.
Deblurring (specifically for motion blur here) involves addressing global motion shifts in the image. As a result, the use of global or multi-scale information is necessary for the deblurring network. Thus, the networks based on the U-shape architecture all perform well on this task. In contrast, SwinIR, which focuses more on local information processing, performs much worse than other networks.
Similar phenomena can be observed for dehazing. Due to the involvement of the depth information in the haze model, the ability to use large-range or even global information is crucial. Besides, dehazing networks are required to handle low-frequency transformations, including alterations in color and contrast, both of which constitute global mappings. Therefore, SwinIR and Uformer, which rely on local spatial information interaction, perform poorly. In contrast, Restormer exhibits exceptional performance.
Deraining is relatively unique in that the rain is unevenly distributed in images, with significant differences between different raindrops and streaks. Thus, there is no clear pattern in the performance of different networks on deraining.
Based on the above results and analysis, we can infer that the acceptable performance of a backbone network on a specific task is predicated on meeting the functional requirements of that task. It is notable that Restormer obtains exceptional task generality. This can be attributed to several factors: 1) The U-shape architecture enables the network to accommodate large-size input. 2) The transposed self-attention allows direct interaction of global information. 3) The presence of depth-wise convolution enables the network to process spatial information effectively. In summary, due to Restormer’s comprehensive functionality, it is capable of meeting the requirements of various tasks.
Figure 5: The network structure of X-Restormer. To enhance the spatial mapping ability of Restormer and create a more general network, we replace half of the transposed self-attention blocks in Restormer with spatial self-attention blocks. For TSA, we retain the preliminary multi-Dconv head transposed attention (MDTA) used in Restormer. For SSA, we adopt the overlapping cross-attention (OCA) in HAT [8].
4 General Backbone Network Design
Based on the above benchmark experiment, we believe that the principle of designing a general backbone network should be to ensure that the network can fulfill the functional requirements of all tasks. As Restormer shows relatively good task generality compared to the other networks, we select it as the starting point to design a more general network. By pinpointing and addressing the limitation of Restormer, we present an initial version of a general image restoration backbone network in this section.
Limitation of Restormer.
In the benchmark experiments, Restormer shows inferior performance to SwinIR on SR, particularly on Urban100. The qualitative comparison also indicates this phenomenon in Figure 4. Although Restormer can directly utilize global information, it fails to reconstruct fine textures even for self-repeated patterns, while SwinIR is capable of doing so. This discrepancy can be partially attributed to the U-shape architecture employed by Restormer. However, this architecture is integral to certain tasks. Another contributing factor is that Restormer relies on depth-wise convolution for spatial information interaction, and its spatial mapping capability is relatively weaker than spatial self-attention in SwinIR. To overcome this limitation, we incorporate a spatial self-attention module into Restormer to design a more powerful backbone network.
Network structure.
In Figure 5, we present our proposed backbone network, denoted as X-Restormer. We choose the U-shape architecture to build the network. In contrast to Restormer, we replace half of the transposed self-attention blocks (TSAB) with spatial self-attention blocks (SSAB) to enhance the ability of spatial information interaction. Given an input feature $F_{in}$, the two blocks process it alternately as:
$F_{t}=F_{in}+TSA(LN(F_{in})),$ (6)
$F_{t\underline{~{}}out}=F_{t}+FFN(LN(F_{t})),$ (7)
$F_{s}=F_{t\underline{~{}}out}+SSA(LN(F_{t\underline{~{}}out})),$ (8)
$F_{out}=F_{s}+FFN(LN(F_{s})),$ (9)
where $F_{t}$, $F_{t\underline{~{}}out}$, $F_{s}$ and $F_{s\underline{~{}}out}$ represent the intermediate feature in TSAB, the output of TSAB, the intermediate feature in SSAB and the output of SSAB. $F_{out}$ means the output of the two consecutive blocks, and also serves as the input for the following two blocks. $TSA(\cdot)$ and $SSA(\cdot)$ indicate transposed self-attention (TSA) and spatial self-attention (SSA) modules. $LN(\cdot)$ denotes layer normalization and $FFN(\cdot)$ represents the feed-forward network.
Table 3: Quantitative results on ×$\times 4$ image SR. * means the SwinIR model is pre-trained on ×$\times 2$ SR.
Table 4: Quantitative results on image denoising with the Gaussian noise level $\sigma=50$.
Model Set5 Set14 BSD100 Urban100 Manga109
RCAN 32.63/0.9002 28.87/0.7889 27.77/0.7436 26.82/0.8087 31.22/0.9173
RCAN-it 32.69/0.9007 28.99/0.7922 27.87/0.7459 27.16/0.8168 31.78/0.9217
SwinIR* 32.92/0.9044 29.09/0.7950 27.92/0.7489 27.45/0.8254 32.03/0.9260
IPT 32.64/- 29.01/- 27.82/- 27.26/- -/-
EDT 32.82/0.9031 29.09/0.7939 27.91/0.7483 27.46/0.8246 32.05/0.9254
NAFNet 32.79/0.9010 29.03/0.7919 27.86/0.7463 27.00/0.8112 31.77/0.9216
SwinIR 32.88/0.9041 29.07/0.7944 27.93/0.7490 27.47/0.8258 31.96/0.9255
Restormer 32.94/0.9039 29.06/0.7934 27.91/0.7482 27.32/0.8199 31.96/0.9244
X-Restormer 33.16/0.9058 29.17/0.7963 28.00/0.7512 27.66/0.8291 32.38/0.9279
Model CBSD68 Kodak24 McMaster Urban100
FFDNet 27.96/- 28.98/- 29.18/- 28.05/-
RNAN 28.27/- 29.58/- 29.72/- 29.08/-
RDN 28.31/- 29.66/- -/- 29.38/-
IPT 28.39/- 29.64/- 29.98/- 29.71/-
DRUNet 28.51/- 29.86/- 30.08/- 29.61/-
SwinIR 28.56/0.8118 29.95/0.8221 30.20/0.8489 29.88/0.8861
Uformer 28.55/0.8130 29.97/ 0.8244 30.16/0.8485 29.98/ 0.8900
Restormer 28.60/0.8130 30.01/0.8237 30.30/0.8517 30.02/0.8898
X-Restormer 28.63/0.8138 30.05/0.8245 30.33/0.8518 30.24/0.8928
Table 4: Quantitative results on image denoising with the Gaussian noise level $\sigma=50$.
Table 5: Quantitative results on motion deblurring.
Table 6: Quantitative results on image deraining.
Model GoPro HIDE RealBlur-R RealBlur-J
SPAIR 32.06/0.953 30.29/0.931 -/- 28.81/0.875
MIMO-UNet+ 32.45/0.957 29.99/0.930 35.54/0.947 27.63/0.837
IPT 32.52/- -/- -/- -/-
MPRNet 32.66/0.959 30.96/0.939 35.99/0.952 28.70/0.873
Uformer 33.05/0.942 30.89/0.920 36.19/0.956 29.09/0.886
NAFNet 33.08/0.942 31.22/0.924 35.97/0.952 28.32/0.857
Restormer 32.92/0.940 31.22/0.923 36.19/0.957 28.96/0.879
X-Restormer 33.44/0.946 31.76/0.930 36.27/0.958 28.87/0.878
Model Test100 Rain100H Rain100L Test1200 Test2800
PreNet 24.81/0.851 26.77/0.858 32.44/0.950 31.36/ 0.911 31.75/0.916
MSPFN 27.50/0.876 28.66/0.860 32.40/0.933 32.39/0.916 32.82/0.930
MPRNet 30.27/0.897 30.41/0.890 36.40/0.965 32.91/ 0.916 33.64/0.938
SPAIR 30.35/0.909 30.95/0.892 36.93/0.969 33.04/0.922 33.34/0.936
SwinIR 30.05/0.900 30.45/0.895 37.00/0.969 30.49/0.893 33.63/0.940
NAFNet 30.33/0.910 32.83/0.914 36.96/0.971 32.58/0.922 32.15/0.933
Restormer 32.03/0.924 31.48/0.905 39.08/ 0.979 33.22/ 0.927 34.21/0.945
X-Restormer 32.21/0.927 32.09/0.914 39.10/0.978 32.31/0.919 33.93/0.945
Table 6: Quantitative results on image deraining.
Table 7: Quantitative results on image dehazing.
Model PFDN FFA-Net AECR-Ne MAXIM DehazeFormer MPRNet NAFNet Restormer X-Restormer
SOTS Indoor 32.68/0.976 36.39/0.989 37.17/0.990 39.72/- 40.05/0.996 40.34/0.994 38.97/0.994 41.97/0.994 42.90/0.995
Specifically, we adopt the Multi-Dconv Transpose Attention (MDTA) as the TSA module. It first generates query (Q), key (K) and value (V) by applying $1\times 1$ convolutions followed by $3\times 3$ depth-wise convolutions. Then, the channel attention matrix of size $\mathbb{R}^{C\times C}$ is calculated by the dot-product of reshaped Q and K followed by a Softmax function. The schematic of TSA is shown in Figure 3. Finally, the result is generated by the dot-product of the attention matrix and V. For SSA, we adopt the Overlapping Cross-Attention (OCA) introduced in the HAT model [8]. We choose this approach because the shifted window mechanism is not suitable for our consecutive blocks, and HAT demonstrates the effectiveness of OCA. For the calculation of OCA, Q is produced by partitioning the input into non-overlapping windows, while K and V are generated by partitioning the input into overlapping windows with a manually set overlapping size. Apart from the different window partition methods, the calculation of OCA is essentially identical to that of standard window self-attention. For FFN, we employ the Gated-Dconv Feed-forward Network (GDFN) architecture, as used in Restormer. Instead of using two $1\times 1$ convolutions to construct an MLP, GDFN first processes input features through two $3\times 3$ depth-wise convolutions and $1\times 1$ convolutions. Then, the resulting features are combined via element-wise multiplication and pass through another $1\times 1$ convolution to produce the final output.
We have also tried multiple design choices for SSAB and TSAB, and the related experiments can be referred to Supp. We emphasize that our design of X-Restormer is not to develop novel architectures or modules to improve the performance on certain tasks, but to enhance the task generality of the network according to the principle of general backbone network design through as simple means as possible.
5 Experiments
5.1 Experimental Setup
We conduct experiments of the proposed X-Restormer on the same datasets used in the benchmark experiment. For the network implementation, the network employs a 4-level encoder-decoder with three times down-sampling and up-sampling. To maintain a similar number of parameters as Restormer, from level-1 to level-4, the numbers of consecutive blocks (containing a TSAB and a SSAB) are [2, 4, 4, 4], attention heads in TSA and SSA are both [1, 2, 4, 8], and channel numbers are [48, 96, 192, 384]. The channel expansion factor in GDFN is 2.66. The overall parameters are 26.06M, slightly less than Restormer of 26.13M. We adopt the same training settings as Restormer in the benchmark experiment to optimize the model. We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, utilizing an initial learning rate of $3e^{-4}$. The learning rate decay follows a cosine scheduler with intervals at 92k and 208k iterations, and the total training iterations are 300K. The input patch size is set to $256\times 256$ and the batch size is set to 32. For data augmentation, we use horizontal and vertical flips. We utilize the $L1$ loss function to train the model. Notably, we do not adopt any training tricks (e.g., ×$\times 2$ SR pretraining or EMA strategy) or testing tricks (e.g., TLC [11]).
Table 8: Quantitative results on All-in-One restoration.
Model SR Denoising Deblurring Deraining Dehazing
×$\times 2$$\times 3$$\times 4$ $\sigma=15$ $\sigma=25$ $\sigma=50$
MPRNet 33.68/0.9301 27.92/0.8295 28.17/0.8043 34.27/0.9280 31.82/0.8901 28.60/0.8119 30.00/0.8812 31.20/0.9068 35.06/0.9874
SwinIR 33.83/0.9301 27.74/0.8351 28.14/0.8043 34.27/0.9283 31.83/0.8906 28.59/0.8143 29.06/0.8519 30.03/0.8983 31.48/0.9823
Uformer 29.99/0.8805 25.02/0.7622 27.88/0.7949 33.86/0.9254 31.42/0.8863 27.87/0.7891 29.64/0.8725 27.53/0.8569 29.92/0.9714
Restormer 34.51/0.9341 23.96/0.7410 28.70/0.8179 34.43/0.9303 32.02/0.8942 28.87/0.8222 30.54/0.8902 31.91/0.9134 36.95/0.9897
NAFNet 34.12/0.9314 23.28/0.7170 28.17/0.8087 34.18/0.9281 31.76/0.8908 28.64/0.8187 30.38/0.8911 31.56/0.9149 30.84/0.9797
X-Restormer 34.72/0.9360 28.04/0.8303 28.81/0.8217 34.67/0.9330 32.26/0.8983 29.12/0.8293 30.85/0.8983 32.27/0.9229 38.24/0.9914
5.2 Experimental Results
We compare our X-Restormer with the top three models in the benchmark experiments (based on the same test configurations) as well as other state-of-the-art approaches (based on the reported performance in their papers) in this section. PSNR/SSIM is provided in following tables. The best and second-best performance results are in bold and underline.
Image SR.
In Table 4, we present the quantitative results of ×$\times 4$ SR on five benchmark datasets: Set5 [2], Set14 [51], BSD100 [31], Urban100 [19] and Manga109 [33]. The state-of-the-art approaches, including RCAN [59], RCAN-it [27], SwinIR [35], IPT [4] and EDT [22] are compared in this experiment. X-Restormer significantly outperforms Restormer by 0.22dB on Set5, 0.34dB on Urban100 and 0.42dB on Manga109. This demonstrates the effectiveness of our design in enhancing the spatial mapping ability of Restormer. Furthermore, X-Restormer surpasses the SOTA method EDT by 0.2dB on Urban100 and 0.35dB on Manga109, indicating the effectiveness of X-Restormer on SR. Despite this, we point out that our method still cannot beat the most powerful SR approaches, such as HAT. We believe that in terms of the SR task, the plain residual in residual architecture designed for SR is still more effective.
Image denoising.
In Table 4, we provide the quantitative results of Gaussian denoising with the noise level $\sigma=50$ on four benchmark datasets: CBSD68 [32], Kodak24 [14], McMaster [58] and Urban100 [19]. The state-of-the-art methods: FFDNet [56], RNAN [61], RDN [62], IPT [4] and DRUNet [57] are compared in this experiment. As one can see that X-Restormer achieves the state-of-the-art performance, surpassing SwinIR by 0.36dB and outperforming Restormer by 0.22dB on Urban100. This demonstrates the superiority of X-Restormer on image denoising.
Image deblurring.
In Table 6, we compare the quantitative results of X-Restormer with the state-of-the-art methods: SPAIR [36], MIMO-UNet+ [10], IPT [4] and MPRNet [49] on both synthetic datasets (Gopro [34] and HIDE [39]) and real-world datasets (RealBlur-R and RealBlur-J [38]). X-Restormer achieves large performance gains over the other models on synthetic datasets, with an improvement of 0.36dB on Gopro compared to NAFNet and 0.54dB on HIDE compared to Restormer. Besides, our X-Restormer obtains the state-of-the-art performance on RealBlur-R and considerable performance on RealBlur-J, showing the effectiveness of our method on real-world motion deblurring scenarios.
Image deraining.
In Table 6, we present the quantitative results of deraining on benchmark datasets Test100 [18], Rain100L [47], Rain100H [47], Test1200 [52] and Test2800[16]. The state-of-the-art methods: PreNet[37], MSPFN [20], MPRNet [49] and SPAIR [36] are compared. X-Restormer outperforms the other models on Test100, Rain100H and Rain100L but performs inferior to Restormer on Test1200 and Test2800. This discrepancy is caused by the variations in degradation produced by different rain models. Nonetheless, our X-Restormer model exhibits comparable performance to state-of-the-art methods, demonstrating its effectiveness on image deraining.
Image dehazing.
In Table 7, we provide the quantitative results on the benchmark SOTS Indoor [21]. We compare the state-of-the-art approaches: PFDN [13], FFA-Net [46], AECR-Net [45], MAXIM [41] and DehazeFormer [40] in this experiment. Notably, our X-Restormer model significantly outperforms Restormer by a large margin of 0.93dB. When compared to the state-of-the-art dehazing method DehazeFormer, our X-Restormer achieves a breakthrough performance gain of 2.85 dB. These results demonstrate the superiority of our proposed method for image dehazing.
All-in-One Restoration.
We conduct experiments on an all-in-one restoration setting, training different networks on five tasks with varying degradation levels (i.e., three SR upscaling factors and $\sigma\in(0,50)$ random Gaussian noises) simultaneously. The sampling probability for each task is the same and the average performance of benchmark datasets is calculated. As shown in Table 8, X-Restormer demonstrates significant advantages in handling multiple tasks concurrently, with its performance far exceeding other methods on all tasks. This indicates that a general backbone network is of great significance for building a general model that can handle multiple image restoration tasks, which can effectively mitigate task conflicts with the performance drops.
Summary.
With enhanced spatial mapping capability, our X-Restormer achieves significant performance improvement compared to Restormer. Specifically, X-Restormer obtains large performance gains against Restormer of 0.42dB (Manga109), 0.22dB (Urban100), 0.54dB (HIDE), 0.61dB (Rain100H) and 0.93dB (SOTS Indoor) on image SR, denoising, deblurring, deraining and dehazing, showing the effectiveness of our design. Moreover, on all these five tasks, X-Restormer obtains state-of-the-art performance. Despite its simplicity, X-Restormer clearly has better task generality than other methods. Furthermore, we demonstrate that a powerful general backbone network has the potential to better handle multiple restoration tasks simultaneously. We hope it can inspire more works on the general image restoration backbone network design.
6 Conclusion
In this paper, we conduct a comparative study of existing image restoration backbone networks to design a general backbone network. Five representative networks are chosen for the benchmark experiment across selected five tasks. The results indicate that comprehensive functionality is crucial for designing a general restoration backbone network. We select Restormer as the baseline and introduce spatial self-attention into it to enhance the spatial information interaction capability. Experimental results show that the designed X-Restormer achieves significant performance improvement and presents the best task generality.
[…]
Supplementary Material
Xiangyu Chen1,2,3 * Zheyuan Li3 * Yuandong Pu2,4 * Yihao Liu2 Jiantao Zhou1 Yu Qiao2,3 Chao Dong2,3
1University of Macau 2Shanghai AI Lab 3Shenzhen Institutes of Advanced Technology, CAS 4Shanghai Jiao Tong University
https://github.com/Andrew0613/X-Restormer
2 Alleviating optimization conflicts
In the manuscript, we show that our X-Restormer obtains the best average performance on an all-in-one restoration setting. We provide more explanations in this section to illustrate a more general backbone network can effectively alleviate the optimization conflict for multi-task restoration. In Figure 6, we provide the validation curves of different models on image SR×$\times 4$ and deblurring. Our X-Restormer can consistently outperform other networks during training. Restormer with good task generality also performs the second best. However, SwinIR performs poorly on both SR and deblurring. We believe this is because when a network has difficulty optimizing certain tasks, the larger gradients it generates may cause the network to focus on these more challenging tasks. Conversely, a general backbone network can effectively circumvent such optimization conflicts.
Figure 6: Validation curves for SR×$\times 4$ and deblurring.
3 Architecture Schematic Diagrams
In the realm of image restoration networks, prevailing architectures fall into three distinct categories: (a) plain residual-in-residual, (b) U-shaped encoder-decoder and (c) multi-stage progressive architectures. In the section 3.2 of the manuscript, we detail the characteristics of these three kinds of architectures. In Figure 7, we provide the schematic diagrams of these architectures for better understanding.
Figure 7: Architecture illustration for Plain Residual-in-Residual, U-shape and Multi-stage.
4 Complete Benchmark Results
We provide the complete the benchmark results on image SR, denoising, deblurring, deraining and dehazing, as shown in Table 9, Table 10, Table 11, Table 12 and Table 13. Since different methods may not use exactly the same data and calculation ways in their papers, we uniformly calculate the performance using BasicSR toolbox [43] on the totally the same test data for fair comparison.
5 Visual Comparison
We present the visual results of the benchmark experiments with X-Restormer on image SR, denoising, deblurring, deraining and dehazing, as depicted in Figure 8, Figure 9, Figure 10, Figure 11, Figure 12, respectively. Our X-Restormer obtains the best visual quality compared to other networks.
Table 9: Complete benchmark results on ×$\times 4$ image SR. The best and second-best performance results are in bold and underline.
Model Set5 Set14 BSD100 Urban100 Manga109
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
MPRNet 32.57/0.8996 28.90/0.7889 27.78/0.7425 26.88/0.8081 31.44/0.9182
SwinIR 32.88/0.9041 29.07/0.7944 27.93/0.7490 27.47/0.8258 31.96/0.9255
Uformer 30.25/0.8665 27.14/0.7398 27.67/0.7475 25.60/0.7651 31.69/0.9233
NAFNet 32.79/0.9010 29.03/0.7919 27.86/0.7463 27.00/0.8112 31.77/0.9216
Restormer 32.94/0.9039 29.06/0.7934 27.91/0.7482 27.32/0.8199 31.96/0.9244
X-Restormer 33.16/0.9058 29.17/0.7963 28.00/0.7512 27.66/0.8291 32.38/0.9279
Table 10: Complete benchmark results on image denoising with the Gaussian noise level $\sigma=50$. The best and second-best performance results are in bold and underline.
Model CBSD68 Kodak24 McMaster Urban100
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
MPRNet 28.48/0.8087 29.86/0.8193 30.04/0.8447 29.71/0.8847
SwinIR 28.56/0.8118 29.95/0.8221 30.20/0.8489 29.88/0.8861
Uformer 28.55/0.8130 29.97/ 0.8244 30.16/0.8485 29.98/ 0.8900
NAFNet 28.52/0.8098 29.90/0.8204 30.07/0.8455 29.65/0.8840
Restormer 28.60/0.8130 30.01/0.8237 30.30/0.8517 30.02/0.8898
X-Restormer 28.63/0.8138 30.05/0.8245 30.33/0.8518 30.24/0.8928
Table 11: Complete benchmark results on motion deblurring. The best and second-best performance results are in bold and underline.
Model GoPro HIDE RealBlur-R RealBlur-J
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
MPRNet 32.66/0.939 30.96/0.919 35.99/0.952 28.70/0.873
SwinIR 31.66/0.921 29.41/0.896 35.49/0.947 27.55/0.840
Uformer 33.05/0.942 30.89/0.920 36.19/0.956 29.09/0.886
NAFNet 33.08/0.942 31.22/0.924 35.97/0.952 28.32/0.857
SwinIR 31.66/0.921 29.41/0.896 35.49/0.947 27.55/0.840
Restormer 32.92/0.940 31.22/0.923 36.19/0.957 28.96/0.879
X-Restormer 33.44/0.946 31.76/0.930 36.27/0.958 28.87/0.878
Table 12: Complete benchmark results on deraining. The best and second-best performance results are in bold and underline.
Model Test100 Rain100H Rain100L Test1200 Test2800
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
MPRNet 30.29/0.898 30.43/0.891 36.46/0.966 32.94/0.914 33.66/0.939
SwinIR 30.05/0.900 30.45/0.895 37.00/0.969 30.49/0.893 33.63/0.940
Uformer 27.93/0.891 24.06/0.845 35.96/0.965 32.75/0.919 28.22/0.913
NAFNet 30.33/0.910 32.83/0.914 36.96/0.971 32.58/0.922 32.15/0.933
Restormer 32.03/0.924 31.48/0.905 39.08/ 0.979 33.22/ 0.927 34.21/0.945
X-Restormer 32.21/0.927 32.09/0.914 39.10/0.978 32.31/0.919 33.93/0.945
Table 13: Complete benchmark results on dehazing. The best and second-best performance results are in bold and underline.
Model MPRNet SwinIR Uformer NAFNet Restormer X-Restormer
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
SOTS Indoor 40.34/0.994 29.14/0.968 33.58/0.986 38.97/0.993 41.97/0.994 42.90/0.995
Table 14: Ablation study on the connection ways. The better performance results are in bold.
Model Set5 Set14 BSD100 Urban100 Manga109 GoPro HIDE RealBlur-R RealBlur-J
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Parallel 32.61/0.9002 28.92/0.7902 27.78/0.7440 26.85/0.8083 31.36/0.9190 31.53/0.9222 29.67/0.9016 36.12/0.9553 28.95/0.8792
Cascade(ours) 32.82/0.9023 29.05/0.7925 27.85/0.7461 27.17/0.8158 31.77/0.9204 32.25/0.9316 30.50/0.9136 36.24/0.9572 28.92/0.8808
6 Ablation Study
We conduct ablation study to explore the effectiveness of our design for X-Restormer. For fast validation, experiments are implemented on image SR and deblurring tasks with significantly different characteristics, based on a small variant of X-Restormer with the numbers of consecutive blocks of [1,2,2,4], the channel numbers of [36, 72, 144, 288] and the input patch size of $192\times 192$.
Cascade or Parallel.
In Table 14, we show the quantitative results on image SR and deblurring for the cascade and parallel connections of TSAB and SSAB. As one can see that the model using cascade design obtains much better performance on all benchmark datasets. Thus, our X-Restormer uses cascade connection for TSAB and SSAB.
Sequences of TSAB and SSAB.
In Table 15, we present the quantitative results on image SR and deblurring for different sequences of TSAB and SSAB. TSAB-SSAB indicates that the features first go through TSAB and then SSAB. Conversely, SSAB-TSAB represents the opposite process. The two models obtain similar performance on SR, while TSAB-SSAB achieves better performance than SSAB-TSAB on image deblurring. Therefore, we use TSAB-SSAB as the default choice.
Effectiveness of OCA.
In Table 16, we provide the quantitative comparison of standard non-overlapping window self-attention (WSA) with the used overlapping self-attention (OCA) for the choice of spatial self-attention in X-Restormer. The overlapping size is set to 0.5, the same as HAT [8]. We can see that OCA performs better than WSA. Therefore, we adopt OCA in X-Restormer.
Different choices for TSAB and SSAB.
In Table 17, we conduct ablation study on different choices for TSAB and SSAB. MDTA [50] and CAB [59, 8] are two modules that involve channel-wise mapping. Swin [25, 30], Dilation self-attention [53, 42] and OCA [8] are three prevalent options for spatial self-attention. We implement this experiment based on the similar number of parameters and FLOPs for all models. When employing MDTA as the choice for TSAB, the model using OCA performs the best compared to other options for SSAB. When adopting OCA as the choice for SSAB, the model with CAB obtains comparable performance to MDTA+OCA on SR, while it performs inferior on image deblurring. We believe this is because CAB contains many convolutions, which can enhance the spatial mapping ability of the model. However, its capability of global information interaction is weaker than MDTA using channel self-attention. Thus, we use the combination of MDTA and OCA as our default choice.
7 Model Complexity Comparison
In Table 18, we present the model complexity comparison of our X-Restormer with Restormer on image SR and motion deblurring. With similar number of parameters and a relatively larger computational complexity, X-Restormer obtains significant performance improvement over Restormer. We also provide a small variant of X-Restormer, denoted as X-Restormer-s, by reducing the original dimension to 44. With fewer parameters and similar FLOPs (calculated for $256\times 256$ output), X-Restormer-s can also perform much better than Restormer.
Table 15: Ablation study on different sequences of TSAB and SSAB. The better performance results are in bold.
Model Set5 Set14 BSD100 Urban100 Manga109 GoPro HIDE RealBlur-R RealBlur-J
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
SSAB-TSAB 32.82/0.9021 29.05/0.7927 27.86/0.7464 27.19/0.8162 31.80/0.9209 32.13/0.9299 30.45/0.9137 36.13/0.9566 29.01/0.8827
TSAB-SSAB 32.82/0.9023 29.05/0.7925 27.85/0.7461 27.17/0.8158 31.77/0.9204 32.25/0.9316 30.50/0.9136 36.24/0.9572 28.92/0.8808
Table 16: Ablation study on the effectiveness of OCA. The better performance results are in bold.
Model Set5 Set14 BSD100 Urban100 Manga109 GoPro HIDE RealBlur-R RealBlur-J
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
WSA 32.81/0.9021 29.01/0.7914 27.82/0.7453 27.09/0.8138 31.68/0.9203 32.09/0.9295 30.41/0.9129 36.24/0.9570 29.00/0.8822
OCA(ours) 32.82/0.9023 29.05/0.7925 27.85/0.7461 27.17/0.8158 31.77/0.9204 32.25/0.9316 30.50/0.9136 36.24/0.9572 28.92/0.8808
Table 17: Ablation study on different choices for TSAB and SSAB. The best and second-best performance results are in bold and underline.
Model Params(M) FLOPs(G) Set5 Set14 BSD100 Urban100 Manga109 GoPro HIDE RealBlur-R RealBlur-J
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
MDTA+Swin 10.6 77.0 32.79/0.9018 29.00/0.7913 27.82/0.7451 27.06/0.8130 31.67/0.9204 32.01/0.9289 30.32/0.9118 36.23/0.9554 28.96/0.8805
MDTA+Dilation 10.6 76.7 32.74/0.9013 28.98/0.7904 27.81/0.7446 27.00/0.8110 31.58/0.9189 31.85/0.9266 30.22/0.9107 36.19/0.9563 28.90/0.8779
CAB+OCA 10.8 73.6 32.74/0.9016 29.04/0.7927 27.85/0.7460 27.18/0.8165 31.78/0.9221 32.47/0.9341 30.39/0.9117 36.02/0.9546 28.71/0.8739
MDTA+OCA(ours) 10.6 78.7 32.82/0.9023 29.05/0.7925 27.85/0.7461 27.17/0.8158 31.77/0.9204 32.25/0.9316 30.50/0.9136 36.24/0.9572 28.92/0.8808
Table 18: Model Complexity Comparison. FLOPs are calculated for $256\times 256$ output.
Model Params(M) FLOPs(G) Set5 Set14 BSD100 Urban100 Manga109 GoPro HIDE RealBlur-R RealBlur-J
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Restormer 26.1 141.0 32.94/0.9039 29.06/0.7934 27.91/0.7482 27.32/0.8199 31.96/0.9244 32.92/0.9419 31.22/0.9226 36.19/0.9571 28.96/0.8785
X-Restormer-s 22.3 140.9 33.04/0.9047 29.18/0.7963 27.97/0.7505 27.63/0.8287 32.27/0.9256 33.41/0.9452 31.52/0.9277 36.24/0.9572 28.75/0.8730
X-Restormer 26.0 164.3 33.16/0.9058 29.17/0.7963 28.00/0.7512 27.66/0.8291 32.38/0.9279 33.44/0.9459 31.76/0.9299 36.27/0.9581 28.87/0.8780
Figure 8: Visual comparison on ×$\times 4$ image SR.
Figure 9: Visual comparison on image denoising with the Gaussian noise level $\sigma=50$.
Figure 10: Visual comparison on motion deblur.
Figure 11: Visual comparison on image deraining.
Figure 12: Visual comparison on image dehazing.